OpenStudy (anonymous):

Given that f(x) is 2x^2/5x^2-9x-2, a) find the limit of f(x) as x approaches infinity b) state the domain of the function

OpenStudy (psymon):

Well, there is a sort of trick to finding these rational fraction limits fast. But the drawn outway is totake the largest power of x in the denominator and divide every term in the problem by that power of x. So in your case, the largest power in the denominator is x^2. So you would then take x^2 and divide all terms in the problem by x^2.

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

but how would the denominator turn out if you divide by x^2? because i get the 2/5 part but what about the 9x-2?

OpenStudy (psymon):

Right, you would have -9/x and -2/x^2. Well if you took those fractions separately to infinity what would you get?

OpenStudy (anonymous):

I don't know :(

OpenStudy (psymon):

Well let's think of it this way. If I give you: \[\frac{ 1 }{ x }\] now this is an example of when the power of x is higher in the denominator than in the numerator. So let's increasing raise the value of x in this fraction: \[\frac{ 1 }{ 10 }->\frac{ 1 }{ 100 }->\frac{ 1 }{ 1000000000 }\]Can you see what the results are approaching if I were to continue?

OpenStudy (anonymous):

It would get closer and closer to 0?

OpenStudy (psymon):

Correct. So thisis exactly what happens with the -9x and the -2. Because they have higher powers of x in the bottom, they go to 0. So if all of those go to 0, you're only left with 2/5, which is the answer.

OpenStudy (anonymous):

thank you! How do I find the domain then?

OpenStudy (psymon):

The normal way you'd find a domain of a function :P

OpenStudy (anonymous):

ahah (-infinity, 0}?

OpenStudy (psymon):

And then the (0,infinity)

OpenStudy (psymon):

Well, you need to factor the bottom xD

OpenStudy (psymon):

Set the denominator = to 0 and factor it to solve for x.

OpenStudy (psymon):

You know how to factor 5x^2-9x-2, correct?

OpenStudy (anonymous):

lol no. I don't know how to do it when there's a number in front of the x^2

OpenStudy (psymon):

Ah. Yeah, I try and see the difficulty of the problem you're doing and sometimes I make incorrect assumptions based on it xD Okay, when the number in front of x^2 is not 1, we just have extra steps. You still need to find factors of (5)(-2) that add up to -9.

OpenStudy (anonymous):

so -5, -1, -2, and 1? no that doesn't work. umm..

OpenStudy (psymon):

Well, 5 and -2 is -10. Factors of -10 that add up to -9 would be -10 and +1

OpenStudy (anonymous):

so x+10 and x-1 would be the factors. and then set that equal to 0?

OpenStudy (psymon):

No, -10 and + 1. But there's more work than that when we have that number in front of x^2 to not be 1. What we do is replace the middle term with our factors. So we would rewrite out function as: \[5x ^{2}-10x+x-2\] Then we need to do factoring by grouping.

OpenStudy (anonymous):

(x-2) and (x-2)?

OpenStudy (anonymous):

bring the 5x out in front

OpenStudy (anonymous):

so 5x (x-2) and (x-2)?

OpenStudy (psymon):

Well, you have to factor something out of each parenthesis. Its just in casethe only thing you can factor out is a 1. So we would have: 5x(x-2)+1(x-2), which means our factors are (5x+1)(x-2)

OpenStudy (anonymous):

and thenn set it equal to 0!

OpenStudy (psymon):

Each one, yes.

OpenStudy (psymon):

So as you can see, this means we have to exclude -1/5 and 2 from our domain. Do you know how to write that in interval notation?

OpenStudy (anonymous):

(-infinity, -1/5} U {2, infinity} ?

OpenStudy (psymon):

Not quite. you left out an interval :3

OpenStudy (psymon):

We still have to include numbers in between -1/5 and 2. Those numbers still do exist.

OpenStudy (anonymous):

so.. is that an intersection? [1/5, 2]?

OpenStudy (psymon):

We still use parenthesis for that portion.

OpenStudy (psymon):

\[(-\infty,-\frac{ 1 }{ 5 })\cup(-\frac{ 1 }{ 5 },2)\cup(2,\infty)\]

OpenStudy (anonymous):

thank you so much! I actually learned how to do several concepts with your help. Thank you!

OpenStudy (psymon):

Yep, np. If you got other problems to look at while we're here, then go for it xD

OpenStudy (anonymous):

I have one last question.. It's another domain question but more difficult in my opinion.

OpenStudy (psymon):

Go for it.

OpenStudy (anonymous):

Find the domain. Use interval notation. f o g, where f(x) = 1/(x+2) and g(x) = 4/(x-1)

OpenStudy (psymon):

Well, the first thing we need is the domain of f and the domain of g before we do the fog part.

OpenStudy (anonymous):

It is already factored so I can just set each to 0?

OpenStudy (anonymous):

so x = -2 amd x = 1

OpenStudy (psymon):

Right. So we need to exclude those from the domain. Now we do the fog part. Do you have any idea how to go about doing that?

OpenStudy (anonymous):

plug in f into g? somehow..

OpenStudy (psymon):

g into f.

OpenStudy (psymon):

So that means the entire function of g(x) mustgo into every x that is in f(x).

OpenStudy (anonymous):

g(x) = 4/0 though. Isn't that undefined?

OpenStudy (psymon):

It equals 4/(x-1), just like you wrote. Now you plug that into the x in f(x)

OpenStudy (anonymous):

so 1/ (4/x-1) +2

OpenStudy (anonymous):

would that change into 1 (x-1)/ 4+2?

OpenStudy (psymon):

Correct. Now you need to simplify it best you can. Well, I'm guessing they want you to at least.

OpenStudy (psymon):

No, it wouldnt. Not that easy unfortunately x_x

OpenStudy (anonymous):

can i multiply the whole denominator to the top and cancel out the denominator then?

OpenStudy (psymon):

You could do a few things, whichever seems best to do. You could literally do the division ny flipping and multiplying like: \[\frac{ 1 }{ \frac{ 4 }{ x-1 } +2}=\frac{ 1 }{ 1 }*(\frac{ x-1 }{ 4 }+\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

so then x-1/4 + 2/4 = x-1+2/4 which equals x+1/4?

OpenStudy (psymon):

No, looks like you did some simplification wrong. What I would try to do is multiply top and bottom by x-1.

OpenStudy (anonymous):

um so x-1/4 + 2x-2?

OpenStudy (psymon):

Correct. Now just simplify that a slight bit :P

OpenStudy (anonymous):

can i factor out a (x-1) or a 2 anywhere?

OpenStudy (psymon):

Well, the 4 - 2 makes it 2+2x, meaning you can make the bottom 2(1+x)

OpenStudy (anonymous):

then should I multiply the top by 2?

OpenStudy (psymon):

No need, you can leave it as is. Now that we finally have fog, what must we exclude from the domain?

OpenStudy (anonymous):

the -2 and 1 from one of the beginning steps.

OpenStudy (psymon):

Yeah, but what about our new function, fog? We have to find numbers from that to exclude, too

OpenStudy (anonymous):

(- infinity, -2) U (-2, 1) U (1, infinity) ?

OpenStudy (psymon):

You're forgetting fog that we just found xD

OpenStudy (anonymous):

so I plug in the -2 and 1 to the x spots which gives me 3/2 and 0

OpenStudy (psymon):

No, you're overthinking it. Just look ONLY at what we found for fog. We have (x-1)/(2(1+x)) What must we exclude from the domain of that?

OpenStudy (anonymous):

1

OpenStudy (psymon):

Nope, we only look at the denominator.

OpenStudy (anonymous):

the two? or the (1+x)?

OpenStudy (psymon):

The 2 doesnt matter really, only the 1+x part.

OpenStudy (anonymous):

so we have to take that out of the domain. but how?

OpenStudy (psymon):

How did we find the domain at the beginning?

OpenStudy (anonymous):

we factored and then set it equal to 0.

OpenStudy (psymon):

Right. So set 2(x+1) = 0 and get x

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

(- infinity, -1] U [-1, infinity)?

OpenStudy (psymon):

Right, that's what I was trying to get you to find xD So we have to exclude everything we found. So we have to exclude -2, -1, and 1.

OpenStudy (anonymous):

oh great. um (-infinity, -2) U (-2, 0) U [1, infinity}?

OpenStudy (psymon):

All parenthesis, no brackets. And we never excluded 0, we can have 0. This is our interval and see if it makes sense: \[(-\infty,-2)\cup(-2,-1)\cup(-1,1)\cup(1,\infty)\]

OpenStudy (anonymous):

YAY, AWESOME! THANK YOU SO MUCH! Honestly learned much more than I did with my pre-calc teacher last year. Thank you!

OpenStudy (psymon):

Lol, np. Surprised you had that limit problem, though. People learn that in regular calculus xD

OpenStudy (anonymous):

this was our summer packet to enter into AP Calc AB.. sort of a struggle after the whole summer, reallly.

OpenStudy (psymon):

Ah, gotcha. Well, limits are the very first thing you do in calculus, so if you can preview them, thats great. After limits is derivatives. So if you got those to look at it might help you out a ton to go over some of that stuff :P

OpenStudy (anonymous):

Thank you so much again! :)

OpenStudy (psymon):

Yep, np :3